graph the equation using the slope and the y-intercept
y=4/3x+2

Answer by algebrahouse.com(1659)

"graph the equation using the slope and the y-intercept"

Slope-intercept form is y = mx + b
m is the slope {rise/run}
b is the y-intercept {where the line crosses the y-axis}

y = (4/3)x + 2
slope = 4/3
y-intercept = 2

To graph:
- put a point on 2 on the y-axis
- from there, go up 4 and to the right 3 and put another point
{since the slope is 4/3, the line rises 4 for every 3 it goes to the right}
- draw a line through the two points
www.algebrahouse.com